# The Brownian loop measure on Riemann surfaces and applications to length spectra

### v1

@@ 3 Brownian loop measure in a homotopy class
Since the Brownian loop measure is conformally invariant, choosing the unique constant curvature metric (up to scaling) on $X$ allows us to explicitly compute the total mass. If it is not hyperbolic, then $X$ is conformal to one of ^,,$\hat{\mathbb{C}},\mathbb{C},\mathbb{C}\setminus\{0\}$ or flat tori. On $\hat{\mathbb{C}},\mathbb{C}$ and $\mathbb{C}\setminus\{0\}$, every homotopy class of closed curves is either homotopically trivial or homotopic to a puncture point, and hence the mass of Brownian loops is infinite. So, we will only consider complete hyperbolic surfaces and flat tori.
@@ 3.1 On hyperbolic surfaces
Let $X$ be a Riemann surface without boundary endowed with the unique complete hyperbolic metric. Then it is isometric to a quotient of the upper half plane $\𝑋\ΓX=\Gamma\backslash\mathbb{H}$, where $\Gamma$ is a Fuchsian group (i.e., a discrete subgroup of ${\rm PSL}(2,\mathbb{R})$ which acts on $\mathbb{H}$ by $\big{(}\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\big{)}:z\mapsto\frac{az+b}{cz+d}$) without elliptic elements. There exists a unique closed geodesic in every free homotopy class that is not homotopic to a point. Let $\gamma$ be an oriented closed geodesic on $X$, and denote $\ell_{\gamma}(X)=\ell(\gamma)$ be the length of it. We write
@@
Brownian motion on $X$ starting from $x$ is a Markov process that only depends on the local geometry of $X$. Hence, the measure on Brownian paths $\mathbb{W}_{x}^{t}(X)$ lifts to $\mathbb{W}_{z}^{t}(\mathbb{H})$. For a Brownian path $\eta$ in $\mathbb{H}$ to be projected down to a curve in $\mathcal{C}_{X}(\gamma)$, the endpoint of $\eta$ has to be $h\cdot z\in\mathbb{H}$ for some $h\in[\gamma]$. We obtain the identity via the disintegration with respect to the endpoint $y$ of the lift of $\mathbb{W}_{x}^{t}(X)$ and evaluate at $y=x$. ∎
@@
where the error function ${\rm Erf}(x)=\frac{2}{\sqrt{\pi}}\int_{0}^{x}\mathrm{e}^{-t^{2}}\,\mathrm{d}t$, which completes the proof. ∎
@@ 4.1 Puncturing the surface
Let $X$ be a complete hyperbolic surface without boundary (perhaps with cusps or funnels). Let $P_{I}=\{p_{k},\,k\in I\}\subset X$ be a non-empty closed subset consisting of countably or finitely many distinct points, and $X^{\prime}=X\setminus P_{I}$ be the Riemann surface with complex structure induced from $X$. Then $X^{\prime}$ admits a unique complete hyperbolic metric conformal to the complex structure.
@@
We note that since $X^{\prime}$ has additional punctures, we cannot apply Theorem 2.2 directly as the conformal factor $\sigma$ between the metrics on $X$ and $X^{\prime}$ is singular at $P_{I}$. However, for all $x\notin P_{I}$, the measure of paths under $\mathbb{W}_{x\to x}(X)$ passing through any point of $P_{I}$ is zero. See, e.g., [33, Cor. 2.26]. On the other hand, a countable set in $X$ has zero volume, as does the total mass of loops starting from a point in $P_{I}$. By Fubini theorem, we have
@@
where $B(p,r)$ is the ball of radius $r$ centered at $p$ in $X$. For $\varepsilon$ small enough, let $X_{\varepsilon}:=X\setminus\cup_{k}B(p_{k},\varepsilon)$. Theorem 2.2 shows that
@@
Similarly, let $T$ be a flat torus and $T^{\prime}=T\setminus P_{I}$ where $P_{I}=\{p_{k}\}_{k\in I}\subset T$ is a non-empty closed subset consisting of countably or finitely many distinct points. Again, $T^{\prime}$ admits a unique complete hyperbolic metric conformal to the complex structure. Same as Corollary 4.1, we still have $\mu_{T}=\mu_{T^{\prime}}$ and $\mu_{T}\{\eta\,|\,\eta\cap P_{I}\neq\emptyset\}=0$.
@@ 4.2 On surfaces with boundary
As we mentioned at the beginning of Section 3, we may also use Lemma 3.2 to compute the Brownian loop measure on a surface $X$ with a non-empty boundary (where the Brownian loops are killed while hitting the boundary). Let $X^{\prime}=X\setminus\partial X$. Since $X^{\prime}$ cannot be one of ^,,$\hat{\mathbb{C}},\mathbb{C},\mathbb{C}\setminus\{0\}$ or flat torus, it admits a unique complete hyperbolic metric. Then, by Theorem 2.2 and Lemma 3.2, the Brownian loop measure can be computed on $X^{\prime}$ with the complete hyperbolic metric:
@@
It was computed in [23, Prop. 3.9] the total mass of non-contractible Brownian loop in the annulus $\mathbb{A}_{r}$, which gives
@@
For $r$ large enough, $\mathbb{A}_{r}\setminus K$ is doubly connected, and there exists $r^{\prime}$ depending on $r$ such that $\mathbb{A}_{r}\setminus K$ is conformally equivalent to $\mathbb{A}_{r^{\prime}}$. It is not hard to see that $\lim_{r\to\infty}r-r^{\prime}=\log|\psi^{\prime}(0)|$. Lemma 3.2 implies that
@@
which is consistent with [18, Lemma 16].
@@
We have $\theta(K)\geq 0$ thanks to the Grunsky inequality [39, Thm. 4.1, (21)].
@@
$\mu_{\hat{\mathbb{C}}}(\{\text{loops with winding $m$ around }0\text{ intersecting }K\})=\frac{\theta(K)}{2\pi^{2}m^{2}}+\frac{1}{2|m|}.$ (4.6)
@@
We write $\tilde{h}=\iota\circ h\circ\iota$ which maps 00 to 00, where $\iota:z\mapsto 1/z$. Then we have $\theta(K)=-\log|\tilde{h}^{\prime}(0)|-\log|f^{\prime}(0)|$. The left-hand side of (4.6) equals
@@
where in the second last equality, we used Corollary 4.6 for the first term; and (4.5) and the fact that $\iota(\Omega^{*})\setminus(\mathrm{e}^{-R}\mathbb{D})$ is conformally equivalent to $\mathbb{A}_{R+\log|\tilde{h}^{\prime}(0)|+o(1)}$ for the second term. ∎
@@
Assume that $X$ is of finite type (i.e., the Euler characteristic is finite) and that $X$ cannot be obtained by removing finitely many (including zero) points from a closed Riemann surface. Then, the total mass of Brownian loops on $X$ that are not homotopic to a point is finite.
@@
Since $X$ is not conformally equivalent to one of ^,,$\hat{\mathbb{C}},\mathbb{C},\mathbb{C}\setminus\{0\}$ or closed tori, it admits a complete hyperbolic metric. We have $\𝑋\ΓX=\Gamma\backslash\mathbb{H}$ where $\Gamma$ is a Fuchsian group without elliptic elements. Let $\mathcal{P}(X)$ be the set of all primitive, oriented, closed geodesics on $X$. By Lemma 3.2, the total mass of Brownian loops of all homotopy classes that are not homotopic to a point is
@@
Denote $N_{X}(L)=\{\gamma\in\mathcal{P}(X)\mid\ell(\gamma)\leq L\}$. If $X$ is elementary (which means $X$ is conformally equivalent to an annulus $\mathbb{A}_{r}$ in our case), $\mathcal{P}(X)$ contains only two elements. We check immediately that the sum in (4.7) is finite. See Remark 4.5.
@@
where $\delta>0$ is the Hausdorff dimension of the limit set of $\Gamma$. If $\delta>\frac{1}{2}$, then $\delta(1-\delta)$ is the smallest eigenvalue of the Laplacian on $X$ [37, 44, 45, 46]. In our case, $X$ has infinite area and hence has no zero eigenvalue, so $\delta$ is strictly smaller than $1$. Then, the total mass of Brownian loops (4.7) is finite. ∎
@@ 4.3 Zeta-regularized determinant of Laplacian
Let us first recall briefly the definition of the zeta-regularized determinant. The spectrum of the positive Laplace–Beltrami operator $\Delta$ on $X$ is an increasing sequence
@@
The zero eigenvalue $\lambda_{0}$ has multiplicity one, and the corresponding eigenfunctions are the constant functions. The $-\log\det\Delta$ is expected to be $-\sum_{k=1}^{\infty}\log\lambda_{k}$, however, which is not well-defined. The zeta-regularization is defined by taking
$-\log\mathrm{det}_{\zeta}\Delta:=\zeta_{X}^{\prime}(0)$
@@
as $t\to 0$ and $T\to\infty$, where $\upgamma\approx 0.5772$ is the Euler-Mascheroni constant, $\chi(X)=2-2\operatorname{genus}(X)$ is the Euler characteristic of $X$, and $\alpha>0$.
@@
The version of the Selberg trace formula for the heat kernel (see, e.g., [42, 30] [8, Ch. 9] [3, Ch. 5]) says that
@@
And directly from this, Naud [34, Eq. (3)] shows that $-\log\mathrm{det}_{\zeta}\Delta$ can be written as the sum of length spectrum
@@
We note that the integral $\int_{L=0}^{\infty}\frac{1}{\mathrm{e}^{L}-1}\ \mathrm{d}N_{X}(L)$ is formally the total mass of Brownian loops homotopic to a primitive geodesic $\sum_{\gamma\in\mathcal{P}(X)}\mu_{X}(\mathcal{C}(\gamma))$. Subtracting from it $\int_{L=0}^{\infty}\frac{1}{\mathrm{e}^{L}-1}\ \mathrm{d}\tilde{\mathop{\rm Li}}(\mathrm{e}^{L})$ renormalizes the contribution of Brownian loops that are homotopic to a long ($L\gg 1$) primitive geodesic as suggested by (4.8).
@@
By prime geodesic theorem (4.8), $|N_{X}(L)-\widetilde{\mathop{\rm Li}}(\mathrm{e}^{L})|=O_{X}(\mathrm{e}^{(1-\epsilon)L})$ as $L\to\infty$ for some $\epsilon>0$ depending on $X$. Hence, the integral in (4.11) converges by integral by parts. The summation in (4.11) converges also because of (4.8) and Lemma 3.2.

### v2

@@ 3 Brownian loop measure in a homotopy class
Since the Brownian loop measure is conformally invariant, choosing a complete metric with constant curvature on $X$ allows us to explicitly compute the total mass. In particular, $X$ is stochastically complete (i.e., Brownian motion is defined for infinite time, see Appendix A). If $X$ is not hyperbolic, then $X$ is conformal to one of ^,,$\widehat{\mathbb{C}},\mathbb{C},\mathbb{C}\setminus\{0\}$ or flat tori. On $\widehat{\mathbb{C}},\mathbb{C}$ and $\mathbb{C}\setminus\{0\}$, every homotopy class of closed curves is either homotopically trivial or homotopic to a puncture point, and hence the mass of Brownian loops is infinite. So, we will only consider complete hyperbolic surfaces and flat tori.
@@ 3.1 On hyperbolic surfaces
Let $X$ be a Riemann surface without boundary endowed with the unique complete hyperbolic metric. Then it is isometric to a quotient of the upper half plane $\𝑋\ΓX=\Gamma\backslash\mathbb{H}$, where $\Gamma$ is a Fuchsian group (i.e., a discrete subgroup of ${\rm PSL}(2,\mathbb{R})$ which acts on $\mathbb{H}$ by $\big{(}\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\big{)}:z\mapsto\frac{az+b}{cz+d}$) without elliptic elements. There exists a unique closed geodesic in every essential (i.e., non-trivial and not homotopic to a cusp) free homotopy class. Let $\gamma$ be an oriented closed geodesic on $X$, and denote $\ell_{\gamma}(X)=\ell(\gamma)$ be the length of it. We write
@@
Brownian motion on $X$ starting from $x$ is a Markov process that only depends on the local geometry of $X$. Hence, the measure on Brownian paths $\mathbb{W}_{x}^{t}(X)$ lifts to $\mathbb{W}_{z}^{t}(\mathbb{H})$. For a Brownian path $\eta$ in $\mathbb{H}$ to be projected down to a curve in $\mathcal{C}_{X}(\gamma)$, the endpoint of $\eta$ has to be $h\cdot z\in\mathbb{H}$ for some $h\in[\gamma]$. We obtain the identity via the disintegration with respect to the endpoints of the lift of $\mathbb{W}_{x}^{t}(X)$:
$\mathbb{W}_{x\to x}^{t}(X)(\mathcal{C}_{X}(\gamma))$
$=\lim_{\varepsilon\to 0+}\frac{\mathbb{W}_{x}^{t}(X)(\{W\in C^{0}([0,t],X)\,|\,W_{t}\in D_{\varepsilon,X}(x),\exists\eta\in\mathcal{C}_{X}(\gamma),d_{X}^{\operatorname{loop}}(W,\eta)<\varepsilon\})}{\mathrm{vol}(D_{\varepsilon,X}(x))}$
$=\lim_{\varepsilon\to 0+}\sum_{h\in[\gamma]}\frac{\mathbb{W}^{t}_{z}(\mathbb{H})(\{W\in C^{0}([0,t],\mathbb{H})\,|\,W_{t}\in D_{\varepsilon,\mathbb{H}}(h\cdot z)\})}{\mathrm{vol}(D_{\varepsilon,\mathbb{H}}(h\cdot z))}$
$=\sum_{h\in[\gamma]}|\mathbb{W}_{z\to h\cdot z}^{t}(\mathbb{H})|,$
where $D_{\varepsilon,X}(x)$ denotes the ball of radius $\varepsilon$ in $X$ centered at $x$, and we used the fact that for $\varepsilon$ small enough (smaller than the injective radius of $X$), $\mathrm{vol}(D_{\varepsilon,X}(x))=\mathrm{vol}(D_{\varepsilon,\mathbb{H}}(h\cdot z))$ for all $h\in\operatorname{PSL}(2,\mathbb{R})$. ∎
@@
where the error function ${\rm Erf}(x)=\frac{2}{\sqrt{\pi}}\int_{0}^{x}\mathrm{e}^{-t^{2}}\,\mathrm{d}t$, and we use the fact that
$\partial_{t}\left({\rm Erf}\left(\frac{L\pm t}{2\sqrt{t}}\right)\right)=\frac{-L\pm t}{2\sqrt{\pi}\ t^{3/2}}\mathrm{e}^{-t/4}\mathrm{e}^{-L^{2}/4t}\mathrm{e}^{\mp L/2}.$ (3.6)
The proof completed. ∎
@@ 4.1 Puncturing the surface
Let $X$ be a complete hyperbolic surface without boundary (may have cusps or funnels and may be of infinite type). Let $P_{I}=\{p_{k},\,k\in I\}\subset X$ be a non-empty, closed, and polar (namely, in all local charts $(U\subset X,\varphi_{U}:U\to\mathbb{C})$, $\varphi_{U}(P_{I}\cap U)$ has zero logarithmic capacity, which is equivalent to $P_{I}$ not being hit by a Brownian motion almost surely, see [36, Thm. 8.20]) subset. For instance, we may take $P_{I}$ to be a closed subset of $X$ which consists of countably or finitely many distinct points. Let $X^{\prime}=X\setminus P_{I}$ be the Riemann surface with complex structure induced from $X$. Then $X^{\prime}$ admits a unique complete hyperbolic metric compatible with the complex structure.
@@
We note that since $X^{\prime}$ has additional punctures, we cannot apply Theorem 2.2 directly as the conformal factor $\sigma$ between the metrics on $X$ and $X^{\prime}$ is singular at $P_{I}$. However, for all $x\notin P_{I}$, the measure of paths under $\mathbb{W}_{x\to x}(X)$ passing through any point of $P_{I}$ is zero. See, e.g., [36, Cor. 2.26]. On the other hand, a countable set in $X$ has zero volume, as does the total mass of loops starting from a point in $P_{I}$. By Fubini theorem, we have
@@
where $B(p,r)$ is the ball of radius $r$ centered at $p$ in $X$. For $\varepsilon$ small enough, $X_{\varepsilon}:=X\setminus\cup_{k}B(p_{k},\varepsilon)$ is non-empty. Theorem 2.2 shows that
@@
Similarly, let $T$ be a flat torus and $T^{\prime}=T\setminus P_{I}$ where $P_{I}=\{p_{k}\}_{k\in I}\subset T$ is a non-empty, closed, and polar subset. Again, $T^{\prime}$ admits a unique complete hyperbolic metric compatible with the complex structure. Same as Corollary 4.1, we still have $\mu_{T}=\mu_{T^{\prime}}$ and $\mu_{T}\{\eta\,|\,\eta\cap P_{I}\neq\emptyset\}=0$.
@@ 4.2 On surfaces with boundary
As we mentioned at the beginning of Section 3, we may also use Lemma 3.2 to compute the Brownian loop measure on a surface $X$ with a non-empty boundary (where the Brownian loops are killed while hitting the boundary). Let $X^{\prime}=X\setminus\partial X$. Since $X^{\prime}$ cannot be one of ^,,$\widehat{\mathbb{C}},\mathbb{C},\mathbb{C}\setminus\{0\}$ or flat torus, it admits a unique complete hyperbolic metric. Then, by Theorem 2.2 and Lemma 3.2, the Brownian loop measure can be computed on $X^{\prime}$ with the complete hyperbolic metric:
@@
It was computed in [25, Prop. 3.9] the total mass of non-contractible Brownian loop in the annulus $\mathbb{A}_{r}$, which gives
@@
For $r$ large enough, $\mathbb{A}_{r}\setminus K$ is doubly connected, and there exists $r^{\prime}$ depending on $r$ such that $\mathbb{A}_{r}\setminus K$ is conformally equivalent to $\mathbb{A}_{r^{\prime}}$.
Since $\psi$ is conformal, for all $\varepsilon>0$, there exists $r_{0}$ large enough, such that for all $r>r_{0}$, the curve $\psi(\mathrm{e}^{-r}\mathbb{S}^{1})$ is contained in the annulus $\{z\in\mathbb{C}\,|\,|\psi^{\prime}(0)|\mathrm{e}^{-r-\varepsilon}<|z|<|\psi^{\prime}(0)|\mathrm{e}^{-r+\varepsilon}\}$. See Figure 2 for an illustration. As $\mathbb{A}_{r^{\prime}}$ is conformally equivalent to the annulus bounded by $\mathbb{S}^{1}$ and $\psi(\mathrm{e}^{-r}\mathbb{S}^{1})$, we have from the monotonicity of the modulus of annuli for inclusion that $-\log|\psi^{\prime}(0)|+r-\varepsilon<r^{\prime}<-\log|\psi^{\prime}(0)|+r+\varepsilon$. This implies $\lim_{r\to\infty}r-r^{\prime}=\log|\psi^{\prime}(0)|$.
Figure 2: Illustration of the annuli in the proof of Corollary 4.6.
Lemma 3.2 implies that
@@
which is consistent with [20, Lemma 16].
@@
We have $\theta(K)\geq 0$ thanks to the Grunsky inequality [42, Thm. 4.1, (21)].
@@
$\mu_{\widehat{\mathbb{C}}}(\{\text{loops with winding $m$ around }0\text{ intersecting }K\})=\frac{\theta(K)}{2\pi^{2}m^{2}}+\frac{1}{2|m|}.$ (4.6)
@@
We write $\iota:z\mapsto 1/z$ and let $\tilde{h}=\iota\circ h\circ\iota$ which maps $\mathbb{D}$ conformally onto $\iota(\Omega^{*})$ while fixing 00. Then we have $|h^{\prime}(\infty)|=1/|\tilde{h}^{\prime}(0)|$ and $\theta(K)=-\log|\tilde{h}^{\prime}(0)|-\log|f^{\prime}(0)|$. The left-hand side of (4.6) equals
@@
where in the second last equality, we used Corollary 4.6 for the first term; and (4.5) and the fact that $\iota(\Omega^{*})\setminus(\mathrm{e}^{-R}\mathbb{D})$ is conformally equivalent to $\mathbb{A}_{R+\log|\tilde{h}^{\prime}(0)|+o(1)}$ for the second term by the same argument as in Corollary 4.6. ∎
@@
Assume that $X$ is a Riemann surface of finite type (i.e., the Euler characteristic is finite) with infinite hyperbolic area. Then, the total mass of Brownian loops on $X$ that are in essential homotopy classes is finite.
@@
Since $X$ is not conformally equivalent to one of ^,,$\widehat{\mathbb{C}},\mathbb{C},\mathbb{C}\setminus\{0\}$ or closed tori, it admits a complete hyperbolic metric. We have $\𝑋\ΓX=\Gamma\backslash\mathbb{H}$ where $\Gamma$ is a Fuchsian group without elliptic elements. Let $\mathcal{P}(X)$ be the set of all primitive, oriented, closed geodesics on $X$. By Lemma 3.2, the total mass of Brownian loops of all essential homotopy classes is
@@
Denote $N_{X}(L)=\{\gamma\in\mathcal{P}(X)\mid\ell(\gamma)\leq L\}$. If $X$ is elementary (which means $X$ is conformally equivalent to an annulus $\mathbb{A}_{r}$ or the once-punctured disk in our case), $\mathcal{P}(X)$ contains only two elements if $X$ is an annulus and none if $X$ is the once-punctured disk. We check immediately that the sum in (4.7) is finite. See Remark 4.5.
@@
where $\delta>0$ is the Hausdorff dimension of the limit set of $\Gamma$. If $\delta>\frac{1}{2}$, then $\delta(1-\delta)$ is the smallest eigenvalue of $-\Delta$ (acting on $L^{2}(X)$) on $X$ [40, 47, 48, 49]. In our case, $X$ has infinite area and hence constant valued functions are not $L^{2}$-eigenfunctions. So it has no zero eigenvalue (since the eigenfunction of zero eigenvalue is of constant value), and $\delta$ is strictly smaller than $1$. Then, the total mass of Brownian loops (4.7) is finite. ∎
@@ 4.3 Zeta-regularized determinant of Laplacian
Let us first recall briefly the definition of the zeta-regularized determinant. The spectrum of the positive Laplace–Beltrami operator $-\Delta$ on $X$ is an increasing sequence
@@
The zero eigenvalue $\lambda_{0}$ has multiplicity one, and the corresponding eigenfunctions are the constant functions. The $\log\det\Delta$ is supposedly [^1] More accurately, we mean $\log\det(-\Delta)$. As only positive operators are considered, we will omit the minus sign before $\Delta$ for shorter notation. $\sum_{k=1}^{\infty}\log\lambda_{k}$, which is not well-defined. We use the zeta-regularization, which is defined by taking
$\log\mathrm{det}_{\zeta}\Delta:=-\zeta_{X}^{\prime}(0)$
@@
as $t\to 0$ and $T\to\infty$, where $\upgamma\approx 0.5772$ is the Euler–Mascheroni constant, $\chi(X)=2-2\operatorname{genus}(X)$ is the Euler characteristic of $X$, and $\alpha>0$.
@@
The version of the Selberg trace formula for the heat kernel (see, e.g., [45] and [8, Ch. 9]) says that
@@
And directly from this, Naud [37, Eq. (3)] shows that $-\log\mathrm{det}_{\zeta}\Delta$ can be written as the sum of length spectrum
@@
Remark 4.12.
The cutoff point $x=2$ of $\widetilde{\mathop{\rm Li}}(x)$ is not special. One can choose another cutoff point, which results in the change of constant $C$.
We note that the integral $\int_{L=0}^{\infty}\frac{1}{\mathrm{e}^{L}-1}\ \mathrm{d}N_{X}(L)$ is formally the total mass of Brownian loops homotopic to a primitive geodesic $\sum_{\gamma\in\mathcal{P}(X)}\mu_{X}(\mathcal{C}(\gamma))$. Subtracting from it $\int_{L=0}^{\infty}\frac{1}{\mathrm{e}^{L}-1}\ \mathrm{d}\widetilde{\mathop{\rm Li}}(\mathrm{e}^{L})$ renormalizes the contribution of Brownian loops that are homotopic to a long ($L\gg 1$) primitive geodesic as suggested by (4.8).
@@
By prime geodesic theorem (4.8), $|N_{X}(L)-\widetilde{\mathop{\rm Li}}(\mathrm{e}^{L})|=O_{X}(\mathrm{e}^{(1-\epsilon)L})$ as $L\to\infty$ for some $\epsilon>0$ depending on $X$. Hence, the integral in (4.11) converges by integration by parts. The summation in (4.11) converges also because of (4.8) and Lemma 3.2.
@@
Appendix A Basics on Brownian motion on Riemannian manifolds
Let us recall first what a standard Brownian motion is and then briefly introduce the Brownian motion on Riemannian manifolds. We only aim to give an intuitive description of them and refer the readers to the more detailed textbook on the topic, e.g., [36, 18].
The Brownian motion in $\mathbb{R}^{n}$ starting from $x=(x_{1},x_{2},\ldots,x_{n})\in\mathbb{R}^{n}$ is a continuous stochastic process $(B_{t})_{t\geq 0}$ in $\mathbb{R}^{n}$, such that the $k$-th coordinate $B^{(k)}_{t}$ a standard one-dimensional Brownian motion starting from $x_{k}$, and all coordinates are independent. In other words, it satisfies
  1. 1.
    for all $k\in\{1,\ldots,n\}$, $B_{0}^{(k)}=x_{k}$,
  2. 2.
    for all $k\in\{1,\ldots,n\}$ and $s<t$, $B_{t}^{(k)}-B_{s}^{(k)}$ has the Gaussian law $\mathcal{N}(0,t-s)$,
  3. 3.
    for all $0=t_{0}<t_{1}<t_{2}<\cdots<t_{l}$, the family of random variables
    $\left\{B_{t_{i}}^{(k)}-B_{t_{i-1}}^{(k)}\,|\,k\in\{1,\ldots,n\},i\in\{1,\ldots l\}\right\}$
    is independent,
  4. 4.
    $t\mapsto B_{t}$ is almost surely continuous.
We will consider the Brownian motion run at speed $2$, that is $W_{t}=B_{2t}$, and consider $W$ as a random variable in the space of $C^{0}([0,\infty),\mathbb{R}^{n})$ endowed with the topology of uniform convergence on compact subsets of $[0,\infty)$. The advantage of considering the speed $2$ Brownian motion is that its diffusion generator is given by the Laplacian $\Delta=\sum_{k=1}^{n}\partial_{kk}$ (instead of $\Delta/2$). In fact, it is straightforward to check that the distribution of $W_{t}$ starting from $x\in\mathbb{R}^{n}$ is given by
$\mathbb{P}_{x}[W_{t}\in\mathrm{d}y]=p(x,y;t)\,\mathrm{d}y=\frac{1}{(4\pi t)^{n/2}}\exp\left(\frac{-\lVert y-x\rVert^{2}}{4t}\right)\prod_{k=1}^{n}\mathrm{d}y_{k}$
where $p(\cdot,\cdot;t)=\mathrm{e}^{t\Delta}$ is the heat kernel at time $t$ and $(p(\cdot,\cdot;t))_{t\geq 0}$ is a semigroup. Moreover, from the definition of Brownian motion, the process $W$ is Markov and satisfies that the conditional law of $W_{t+s}$ given $W|_{[0,s]}$ is $p(W_{s},y;t)\,\mathrm{d}y$. So we also call $p(\cdot,\cdot;t)$ the transition kernel of Brownian motion.
From this, one may define the Brownian motion on $\mathbb{R}^{n}$ directly from the diffusion generator $\Delta$. More precisely, we determine from $\Delta$ the semigroup of heat kernels $(p(\cdot,\cdot;t))_{t\geq 0}$. This allows us to define the finite dimensional marginals, namely the joint law of $(W_{0},W_{t_{1}},\ldots,W_{t_{l}})$, using the heat kernel as the transition kernel. By the semigroup property of the heat kernel, we refine the sequence $0<t_{1}<\cdots<t_{l}$ and define the law of $W$ on a countable dense set of times $I$. Then Kolmogorov extension theorem shows that this construction has a continuous modification, that is, a probability distribution on $C^{0}([0,\infty),\mathbb{R}^{n})$ which has the same finite dimensional marginals when restricted to any finite subset of $I$.
This viewpoint allows us to define the Brownian motion $W$ run at speed $2$ on a Riemannian manifold $(X,g)$ as the diffusion generated by the Laplace–Beltrami operator $\Delta_{g}$ where
$\Delta_{g}(f)=\operatorname{div}_{g}\operatorname{grad}_{g}(f)=\frac{1}{\sqrt{|g|}}\sum_{i,j=1}^{n}\partial_{i}(\sqrt{|g|}g^{ij}\partial_{j}f),$
using the associated heat kernel semigroup $(p_{X,g}(\cdot,\cdot;t))_{t\geq 0}$ as the transition kernel. The book [13] by Chavel contains an extensive discussion on the heat kernel on a Riemannian manifold. Moreover, if $(X,g)$ is complete and whose Ricci curvature is bounded from below, then the Brownian motion is defined for all time, namely, $\int p_{X,g}(x,y;t)\,\mathrm{d}\mathrm{vol}(y)=1$ for all $x\in X$ and $t\geq 0$ (in this case, we say that $(X,g)$ is stochastically complete). See [18, Sec. 4].
We will restrict ourselves to the case of surfaces for which the description of Brownian motion can be simplified and made very concrete, thanks to the fact that if in a local coordinate $g=\mathrm{e}^{2\sigma}(\mathrm{d}x_{1}^{2}+\mathrm{d}x_{2}^{2})$, then $\Delta_{g}=\mathrm{e}^{-2\sigma}(\partial_{11}+\partial_{22})$ (this relation only holds in two-dimension). This implies that if $\sigma\in C^{\infty}(\Omega,\mathbb{R})$ such that $\Omega\subset\mathbb{R}^{2}$ is a domain and $g=\mathrm{e}^{2\sigma}(\mathrm{d}x_{1}^{2}+\mathrm{d}x_{2}^{2})$ is a complete metric on $\Omega$ with bounded curvature, the standard Brownian motion $W$ on $\mathbb{R}^{2}$ run at speed $2$ and the Brownian motion $\widehat{W}$ on $(\Omega,g)$ run at speed $2$, both starting from $x\in\Omega$ and defined for infinite time, can be coupled by a time-change.
More precisely, we let $\mathbb{W}_{x}$ be the law of $W$ (a probability measure on $C^{0}([0,\infty),\mathbb{R}^{2})$), define for each realization of $W$ and $t\geq 0$,
$\widehat{W}_{t}:=W_{s}\quad\text{where}\quad s=\inf\left\{r\geq 0\,\Big{|}\,\int_{0}^{r}\mathrm{e}^{2\sigma(W_{l})}\mathrm{d}l\geq t\right\}.$
We obtain a continuous function $\widehat{W}:\mathbb{R}_{+}\to\Omega$ which is a time-change of $W$ (in fact, $W$ stopped at $\partial\Omega$ suffices). The pushforward measure of $\mathbb{W}_{x}$ by the map $W\mapsto\widehat{W}$, is the law of the Brownian motion on $(\Omega,g)$ starting from $x$.
The Brownian motion on a more general complete Riemannian surface $(X,g)$ is the projection of a Brownian motion by a locally isometric universal covering map $\tilde{X}\to X$, where $\tilde{X}=\mathbb{H}$ or $\mathbb{C}=\mathbb{R}^{2}$ or $\mathbb{S}^{2}$, equipped with a conformal metric $g=\mathrm{e}^{2\sigma}(\mathrm{d}x_{1}^{2}+\mathrm{d}x_{2}^{2})$.
If $X^{\prime}\subset X$ is a subsurface of $(X,g)$, then the law of Brownian motion starting from $x\in X^{\prime}$ killed at $\partial X^{\prime}$ is the probability measure on the space of continuous functions induced from the Brownian motion $\mathbb{W}_{x}$ on $X$ by restricting the continuous path up to the first time exiting $X^{\prime}$. This can also be obtained using the heat kernel with Dirichlet boundary condition on $X^{\prime}$ as the transition kernel [18, Prop. 4.1.3].